Hallux abducto valgus (call it HAV) is a deformation of the big toe that often requires surgery. Doctors used X-rays to measure the angle (in degrees) of deformity in 38 consecutive patients under the age of 21 who came to a medical center for surgery to correct HAV. The angle is a measure of the seriousness of the deformity. Here are the data. 29 23 15 31 15 28 25 17 28 35 21 18 38 22 49 25 16 23 26 31 27 16 23 26 28 21 29 25 22 39 27 22 31 11 16 22 21 26 what is the standard deviation?
Answer #1

Solution: The formula for calculating Standard deviation is given below:

Standard deviation =\sqrt{\frac{\sum (x_{i}-\bar{x})^{2}}{n-1}}

Where x_{i} is the angle if ith patient

\bar{x}=\frac{\sum x_{i}}{n} is mean angle of 38 patients

  n is number of patients

Now lets calculate the mean angle of 38 patients

\bar{x}=\frac{\sum x_{i}}{n}

Angle (x) 23 31 28 17 28 35 21 18 38 49 16 23 26 31 27 16 23 26 28 21 3 39 27 31 16 21 26

i = 947-24.921 38

Now lets find the Standard deviation of the given data:

Angle 16.63781163 3.690443213 9842728532 36.95360111 9842728532 9.479916898 0.006232687 62.74307479 9.479916898 101.5851801 1

Therefore standard deviation is

2066.763-7.4739 Standard deviation 37-7.4739